The college shows that thier student spend an average of $600 in books per year. The standart deviation is $128. If a random sampe of 52 students is surveyed what is the probability that the mena amount spend for books is more than $700
Answer by ewatrrr(24785)

Population: m = 600, sd = 128
Sample of 52 students
P(x-bar >700) = P(z > 100/(128/sqrt(52)) = P(z > 5.6337) = 1 - P(z ≤ 5.6337) = basically zero chance